What two-digit positive integer is one more than a multiple of 2, 3,4, 5 and 6 ?
Answer by ankor@dixie-net.com(6693)

What two-digit positive integer is one more than a multiple of 2, 3,4, 5 and 6 ?
:
Primes representing all these numbers: 2, 2, 3, 5,
:
2 * 2 * 3 * 5 = 60 + 1 = 61